Find area of surface obtained by rotating for = 3.x=2 with 25x56 about the x-axis
Answer #1

0 f(x) = y= 35x-2 Area of sortan obtained A = Lanyar = الى A = 27 ( 35 (x-2] ). Jit (dy dx 2 San Cara sta (mely A = Gr X-2 dy느 69 | 4 (x2) + 4 dx 2 오x-2 A 37 4 우 dx 2 6 A= 3M (x+1) as 32 6 3 | (4x+1) ) 4 3 2. 6 | 길 [th 4x+1) 31일 」 ITY (25) 34 (0) (9)